Please help and show all work thank you
Answer #1

It is given that the given vectors are orthogonal.

So their dot product is zero.

1(-4) + 1(2) + 3(1) +(-3)(k)=0

-4 + 2 + 3 - 3k=0

1 - 3k =0

1 = 3k

k = 1/3